What number comes next 1 2 8 48 384
Found 2 solutions by stanbon, rothauserc:
Answer by stanbon(75887)

What number comes next 1 2 8 48 384
-------
1
1*2 = 2
2*4 = 8
8*6 = 48
48*8 = 384
384*10 = 3840
=====================
Cheers,
Stan H.
======================

Answer by rothauserc(4718)

this sequence is defined as
a(1) = 1
a(n) = 2*(n-1)*a(n-1)
so a(6) = 2*(n-1)*a(5) = 2*5*384 = 3840